Paul Swanson has an opportunity to acquire a franchise from The Yogurt Place, Inc., to dispense frozen yogurt products under The Yogurt Place name. Mr. Swanson has assembled the following information relating to the franchise:

  1. A suitable location in a large shopping mall can be rented for $4,300 per month.
  2. Remodeling and necessary equipment would cost $366,000. The equipment would have a 20-year life and a $18,300 salvage value. Straight-line depreciation would be used, and the salvage value would be considered in computing depreciation.
  3. Based on similar outlets elsewhere, Mr. Swanson estimates that sales would total $460,000 per year. Ingredients would cost 20% of sales.
  4. Operating costs would include $86,000 per year for salaries, $5,100 per year for insurance, and $43,000 per year for utilities. In addition, Mr. Swanson would have to pay a commission to The Yogurt Place, Inc., of 12.0% of sales.

Required:

1. Prepare a contribution format income statement that shows the expected net operating income each year from the franchise outlet.

2-a. Compute the simple rate of return promised by the outlet.

2-b. If Mr. Swanson requires a simple rate of return of at least 22%, should he acquire the franchise?

3-a. Compute the payback period on the outlet.

3-b. If Mr. Swanson wants a payback of two years or less, will he acquire the franchise?
Answer #1
1. The Income Statement would be Amount (in $)
Sales              460,000
Variable Expenses
Cost of ingredients (20% of 460000)          92,000
Commission (12% of 460000)          55,200              147,200
Contribution Margin              312,800
Selling & Administrative Expense
Salaries          86,000
Rent          51,600
Depreciation          17,385
Insurance            5,100
Utilities          43,000              203,085
Net Operating Income              109,715
Depreciation calculation
(366000-18300) 347700
Life 20 years 20
Per year depreciation 17385

Answer 2a)

Simple Rate of return = Annual incremental net operating income / Initial investment

   = 109,715/366,000

= 30%

Answer 2 b) Yes, the franchise would be acquire because it will give rate of return in excess of 22%.

Answer 3 a)

Pay back period = Investment required / annual net cash inflow

= 366000/127100

= 2.88 years   

Net operating income + depreciation = Annual cash inflow

109712 + 17385 = $127,100

Answer 3b) According to the pay back computation, the franchise would not be acquired. The 2.88 years payback is greater than 2 years allowed. payback and simple rate of return can give conflicting signals in this example.